# Long-Context Linear System Identification

### v1

@@ 5 Discussion
We now discuss the rates obtained in Section 4 and compare them with previous results obtained for linear dynamical systems. In particular, we comment the “learning-without-mixing” phenomenon, introduced by Simchowitz et al. (2018) for the first-order linear dynamical systems.
@@
$\mathcal{A}^{\star}=\begin{pmatrix}0_{d}&I_{d}&\cdots&0_{d}\\ \vdots&\ddots&\ddots&\vdots\\ 0_{d}&\cdots&0_{d}&I_{d}\\ A_{p}^{\star}&A_{p-1}^{\star}&\cdots&A_{1}^{\star}\end{pmatrix}\,.$ (12)
We have the relation $X_{t+1}^{(n)}=\mathcal{A}^{\star}X_{t}^{(n)}+\Xi_{t}^{(n)}$, reducing the problem to the $p=1$ case by increasing the dimension from $d$ to $pd$. First, brute-force adapting previous results to this case (e.g. Basu et al., 2019; Simchowitz et al., 2018; Sarkar & Rakhlin, 2019) is not possible since these works assume that the noise covariance of the additive noise added at each step ($\Xi_{t}^{(n)}$ here) is the identity matrix, or at least is positive definite. In our case, the noise covariance is the $pd\times pd$ block-diagonal matrix, with $p-1$ blocks equal to $0_{d}$ and the last one to $I_{d}$. The covariance matrix is thus non-invertible, preventing the use of previous works.
@@
Moreover, adapting the techniques developed in the $p=1$ setting (Sarkar & Rakhlin, 2019) which relies on explicit factorization of the OLS estimator is challenging. In the $p>1$ case, the higher-order dynamics complicate the factorization, and the data matrix takes a Toeplitz form, which is more difficult to handle.
@@
Let $\tau_{\rm mix}$ be the mixing time of the Markov chain [^1] If $\tau_{{\rm mix}}$ does not exist, it is set to $T$. $(X_{t}^{(n)})_{t\geqslant 0}$. In the i.i.d. setting (for which $\tau_{\rm mix}=1$), the OLS estimator obtains the optimal rate $\|{\boldsymbol{\hat{A}}}_{\rm OLS}-{\boldsymbol{A}}^{\star}\|_{F}^{2}=\mathcal{O}(pd^{2}/NT)$, since $pd^{2}$ is the dimension of the inputs. With non-i.i.d. but Markovian data, a naive strategy would be to emulate i.i.d.-ness and take only a sample every $\tau_{\rm mix}$ steps of the trajectory to compute the OLS estimator, thus having data that are approximately i.i.d. while dividing the number of samples by $\tau_{\rm mix}$. This naive estimator would yield $\|{\boldsymbol{\hat{A}}}_{\rm naive}-{\boldsymbol{A}}^{\star}\|_{F}^{2}=\tilde{\mathcal{O}}(\tau_{\rm mix}pd^{2}/N(T-\tau_{\rm mix}))$, where the mixing time appears as a cost of non-i.i.d.-ness.
A rate of convergence that exhibits “learning-without-mixing” is a rate of the form $\|{\boldsymbol{\hat{A}}}-{\boldsymbol{A}}^{\star}\|_{F}^{2}\leqslant Cpd^{2}/NT$ where $C\ll\tau_{{\rm mix}}$. Such a rate means that the matrix ${\boldsymbol{A}}^{\star}$ is learnt without paying the cost of non-i.i.d.-ness. For instance, in the $p=1$ case, Simchowitz et al. (2018) learn without mixing as the mixing time of their Markov chain is governed by $1/(1-\rho)$ where $\rho=\|A^{\star}\|_{\rm op}$, while their rate does not worsen as $\rho$ tends to 1—in fact, $\rho\to 1$ actually improves their rates.
@@
In our case, two components contribute to $\tau_{\rm mix}$. The first component is related to the stability or the excitability of the system and scales as $1/(1-\rho)$, where $\rho=\|M_{{\boldsymbol{A}}^{\star}}\|_{\rm op}$. When $\rho\ll 1$, this component has no impact, while $\rho$ tends to 1, the system is less stable and the Markov chain mixes slower. The second component is directly related to the context length $p$ of the process. Regardless of the factor $1/(1-\rho)$ above, the mixing time of our Markov chain is larger than $p$: since noise is added only in the last block in the recursion $X_{t+1}^{(n)}=\mathcal{A}^{\star}X_{t}^{(n)}+\Xi_{t}^{(n)}$, starting from a given state, $p$ iterations at least are needed to eventually forget this given state. The naive learning-with-mixing benchmark rate is thus $\|{\boldsymbol{\hat{A}}}-{\boldsymbol{A}}^{\star}\|_{F}^{2}\leqslant{\max\left(\frac{1}{1-\rho},p\right)pd^{2}/NT}$, where the factor $\max(1/1-\rho,p)$ takes the two above points into consideration.
The bound presented in Theorem 4.1 takes the form $\tilde{\mathcal{O}}(Cpd^{2}/(N(T-p)))$ where $\tilde{\mathcal{O}}$ hides the logarithmic terms. Importantly, the constant $C$ and the logarithmic terms are independent of the mixing related quantity $\max\left(1/1-\rho,p\right)$.
@@
We now explain the behavior of $\kappa$ defined in Definition 3.6. First, by Lemma B.11, we have that $\sigma_{\min}(L_{\star})\geqslant\frac{1}{D+1}$ and, thus, it is sufficient to upper bound
@@
to control $\kappa$. Equation 13 implies that if the noise at step $i$ contributes to step $j$, as measured by $L_{\star}^{(i,j)}$, at a polynomial rate in $\left(j-i\right)$, then $\kappa$ grows at most logarithmically in $T$. For such a $\kappa$, the resulting dependency on $T$ is of order $\ln T$ and mild. Instead, if it is exponential in $\left(j-i\right)$, then $\kappa$ grows linearly in $T$ and the dependency on $T$ cancels out in the rate.
We use the quantity $\zeta(T)$ to define stable, marginally stable and explosive systems:
@@
stable if: $\quad\zeta(T)=\mathcal{O}(\rho^{T})\quad\text{for some}\quad\rho\leqslant 1\,,$
marginally stable if: $\quad\zeta(T)=\mathcal{O}(T^{k})\quad\text{for some}\quad k\in\mathbb{N}^{\star}\,,$
explosive if: $\quad\zeta(T)=\mathcal{O}(\rho^{T})\quad\text{for some}\quad\rho>1\,.$
Definition 5.1 is similar to the notions of stable, marginally stable and explosive systems considered in (Simchowitz et al., 2018; Sarkar & Rakhlin, 2019) for $p=1$. Let $\rho(A^{\star})\coloneqq\lambda_{\max}(A^{\star})$ be the spectral radius of $A^{\star}$ and $V\Lambda V^{-1}$ be the Jordan normal form of $A^{\star}$. Then,
@@
In particular, for stable systems of Simchowitz et al. (2018); Sarkar & Rakhlin (2019) with $\rho<1$, $\zeta(T)=\mathcal{O}(\rho^{T})$. For marginally stable systems of Sarkar & Rakhlin (2019) with $\rho<1+\frac{\gamma}{T}$ with some constant $\gamma>0$, $\rho(A^{\star})^{j-i}\leqslant e^{\gamma}$ and $\zeta(T)=\mathcal{O}(T^{k})$ for some fixed $k$ that depends on the largest Jordan block of $A^{\star}$. For explosive systems of Sarkar & Rakhlin (2019) with $\rho>1$, $\zeta(T)=\mathcal{O}(\rho^{T})$. Thus, Definition 5.1 provides a general categorization of the systems based on the growth of $\zeta(T)$ in $p>1$ case. Furthermore, our analysis yields sharp rates for stable and marginally stable systems previously considered only in the $p=1$ setting.
@@
Our analysis is based on the assumption that the diameter $D$ of the search space is bounded and, hence, not directly applicable to the OLS estimator in Equation 4. However, Corollary B.7 in Section B.1 extends the results of Theorems 4.1, 4.2 and 4.3 to minimizers without a constraint on the diameter of the search space. This extension does not change the rates but requires the additional assumption that $NT=\tilde{\Omega}(p^{\prime 2}dr)$. [^2] We use the convention that $r=d,p^{\prime}=p$ for Theorem 4.1. In the case of Theorem 4.1, this corresponds to a result for the OLS estimator, but necessitating a number of samples quadratic in context length. Below, we comment on why the diameter restrictions is required when $NT\ll p^{\prime 2}dr$.
As mentioned earlier in comparison with (Simchowitz et al., 2018; Sarkar & Rakhlin, 2019), the simple OLS factorization in the $p=1$ case does not generalize to the $p>1$ and the data matrix has a Toeplitz structure that is more difficult to control. In order to deal with these issues, as explained in the sketch of proof in Section 6, we rely on techniques from empirical process theory. These techniques are applied to quantify the probability of the event in Equation 14, which hold for any empirical risk minimizer of the square loss. This leads us to the study of the concentration of the martingales defined in Equation 16 around their predictable variation, which is a key step in our analysis. A uniform concentration is possible only if there is a uniform lower bound on the variations of the martingales, which can be achieved using a set of well-behaved matrices $\|M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}}\|_{F}/\|M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}}\|_{\rm op}$. In order to translate these conditions on the design matrices without additional dimensional dependencies, we introduce the operator norm constraint.
@@
The misspecification result in Theorem 4.3 requires the additional assumption given in Assumption 3.9. In Remark B.4, we show that a good upper bound on $D^{\prime}$ is possible whenever $D<1$, i.e., the system is stable, by using the upper bound $\|L_{\star}\|_{\rm op}\leqslant 1/(1-D)$.
6 Sketch of Proof
We provide a sketch of proof for Theorem 4.1. The proofs of Theorems 4.2 and 4.3 are similar and can be found in Appendix B. In the following, $\Delta_{{\boldsymbol{A}}}$ is a shorthand for $M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}}$.
@@ 6 Sketch of Proof
$\forall{\boldsymbol{A}}\in\mathcal{A}^{\prime}(D)\,,\quad{{\left\|\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right\|}}^{2}\geqslant 2\operatorname{Tr}\left(E^{\top}L_{\star}^{\top}\Delta_{{\boldsymbol{\hat{A}}}}^{\top}E\right)\,,$ (15)
@@
Fix a ${\boldsymbol{A}}\in\mathcal{A}^{\prime}(D)$ and study the martingale series defined through the differences sequences
$d_{t,i}^{(n)}=\left(\left({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}\right)x_{t}^{(n)}\right)_{i}\left(\xi_{t}^{(n)}\right)_{i}/\sigma^{2}\,,$ (16)
@@
$Y_{{\boldsymbol{A}}}=\sum_{n,t,i}d_{t,i}^{(n)}=\sum_{n,t}\left\langle\left({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}\right)x_{t}^{(n)},\xi_{t}^{(n)}\right\rangle=\frac{1}{\sigma^{2}}\operatorname{Tr}\left(E^{\top}L_{\star}^{\top}\Delta_{{\boldsymbol{A}}}^{\top}E\right)\,,$
and the quadratic predictable variation of the series is
$W_{{\boldsymbol{A}}}=\sum_{n,t,i}\mathbb{E}_{\left(\xi_{t}^{(n)}\right)_{i}}\left[\left(d_{t,i}^{(n)}\right)^{2}\right]=\frac{1}{\sigma^{2}}\sum_{n,t}{{\left\|\left({\boldsymbol{A}}^{\star}-{\boldsymbol{A}}\right)x_{t}^{(n)}\right\|}}^{2}=\frac{1}{\sigma^{2}}{{\left\|\Delta_{{\boldsymbol{A}}}L_{\star}E\right\|}}^{2}\,.$
The condition that is asked in Equation 15 is then that the sum of the differences $Y_{{\boldsymbol{A}}}$ is large compared to the quadratic predictable variation $W_{{\boldsymbol{A}}}$:
$\mathcal{E}\coloneqq\left\{\forall{\boldsymbol{A}}\in\mathcal{A}^{\prime}(D):W_{{\boldsymbol{A}}}\leqslant 2Y_{{\boldsymbol{A}}}\right\}\,,$
@@
$\begin{split}\mathbb{P}\left(Y_{\boldsymbol{A}}\geqslant r_{Y}\,,W_{\boldsymbol{A}}^{R}\leqslant r_{W}\right)&\leqslant\exp\left(-\frac{r_{Y}^{2}/2}{r_{W}+Rr_{Y}}\right)\,,\\ &\enspace\text{where}\enspace W_{{\boldsymbol{A}}}^{R}=W_{\boldsymbol{A}}+\sum_{n,t,i}\mathds{1}_{d_{t,i}^{(n)}>R}\left(d_{t,i}^{(n)}\right)^{2}\,,\end{split}$ (17)
and $r_{Y}>0,r_{W}>0$ and $R>0$ are arbitrary constants. As the noise is sub-Gaussian, it is possible to upper bound $W_{{\boldsymbol{A}}}^{R}$ with $W_{{\boldsymbol{A}}}$:

### v2

@@ 5 Discussion
We now discuss the rates obtained in Section 4 and compare them with previous results obtained for linear dynamical systems. In particular, we comment on the “learning-without-mixing” phenomenon, introduced by Simchowitz et al. (2018) for the first-order linear dynamical systems.
@@
$\mathcal{A}^{\star}=\begin{pmatrix}0_{d}&I_{d}&\cdots&0_{d}\\ \vdots&\ddots&\ddots&\vdots\\ 0_{d}&\cdots&0_{d}&I_{d}\\ A_{p}^{\star}&A_{p-1}^{\star}&\cdots&A_{1}^{\star}\end{pmatrix}\,.$
We have the relation $X_{t+1}^{(n)}=\mathcal{A}^{\star}X_{t}^{(n)}+\Xi_{t}^{(n)}$, reducing the problem to the $p=1$ case by increasing the dimension from $d$ to $pd$. First, a brute-force adaptation of previous results to this case (e.g. Basu et al., 2019; Simchowitz et al., 2018; Sarkar & Rakhlin, 2019) is not possible since these works assume that the noise covariance of the additive noise added at each step ($\Xi_{t}^{(n)}$ here) is the identity matrix, or at least is positive definite. In our case, the noise covariance is the $pd\times pd$ block-diagonal matrix, with $p-1$ blocks equal to $0_{d}$ and the last one to $I_{d}$. The covariance matrix is thus non-invertible, preventing the use of previous works.
@@
Moreover, adapting the techniques developed in the $p=1$ setting (Sarkar & Rakhlin, 2019) which rely on explicit factorization of the OLS estimator is challenging. In the $p>1$ case, the higher-order dynamics complicate the factorization, and the data matrix takes a Toeplitz form, which is more difficult to handle.
@@
We explain why our rates exhibit “learning-without-mixing”. We begin by defining “learning-with-mixing” and discussing the factors that influence the mixing time $\tau_{{\rm mix}}$. We then introduce the concept of “learning-without-mixing” as exemplified by Simchowitz et al. (2018) and show that our bounds exhibit similar properties.
@@
Let $\tau_{\rm mix}$ be the mixing time of the Markov chain $(X_{t}^{(n)})_{t\geqslant 0}$. In the i.i.d. setting (for which $\tau_{\rm mix}=1$), the OLS estimator obtains the optimal rate $\|{\boldsymbol{\hat{A}}}_{\rm OLS}-{\boldsymbol{A}}^{\star}\|_{F}^{2}=\mathcal{O}(pd^{2}/NT)$, since $pd^{2}$ is the dimension of the inputs. With non-i.i.d. but Markovian data, a naive strategy would be to emulate i.i.d.-ness and take only a sample every $\tau_{\rm mix}$ steps of the trajectory to compute the OLS estimator, thus having data that are approximately i.i.d. while dividing the number of samples by $\tau_{\rm mix}$. This naive “learning-with-mixing” estimator would yield $\|{\boldsymbol{\hat{A}}}_{\rm naive}-{\boldsymbol{A}}^{\star}\|_{F}^{2}=\tilde{\mathcal{O}}(\tau_{\rm mix}pd^{2}/NT)$, where the mixing time appears as a cost of non-i.i.d.-ness and $\tilde{\mathcal{O}}$ hides the logarithmic terms in problem parameters $p,d,N,T$.
@@
In our case, two components contribute to the mixing time, $\tau_{\rm mix}$. The first component is related to the stability or the excitability of the system and scales as $1/(1-\rho)$, where $\rho=\|M_{{\boldsymbol{A}}^{\star}}\|_{\rm op}<1$. When $\rho\ll 1$, this component has no impact, while $\rho$ tends to 1, the system is less stable and the Markov chain mixes more slowly. The second component is directly related to the context length $p$ of the process. Regardless of the factor $1/(1-\rho)$ above, the mixing time of our Markov chain is larger than $p$: since noise is added only in the last block in the recursion $X_{t+1}^{(n)}=\mathcal{A}^{\star}X_{t}^{(n)}+\Xi_{t}^{(n)}$, starting from a given state, $p$ iterations at least are needed to eventually forget this given state. The naive learning-with-mixing benchmark rate is thus $\|{\boldsymbol{\hat{A}}}-{\boldsymbol{A}}^{\star}\|_{F}^{2}\leqslant{\max\left(1/\left(1-\rho\right),p\right)pd^{2}/NT}$.
In contrast, a rate of convergence that exhibits “learning-without-mixing” is a rate of the form $\|{\boldsymbol{\hat{A}}}-{\boldsymbol{A}}^{\star}\|_{F}^{2}\leqslant Cpd^{2}/NT$ where $C\ll\tau_{{\rm mix}}$. Such a rate means that the matrix ${\boldsymbol{A}}^{\star}$ is learned without paying the cost of non-i.i.d.-ness. For instance, in the $p=1$ case, the rate of Simchowitz et al. (2018) does not worsen as $\rho$ tends to 1—in fact, $\rho\to 1$ actually improves their rates.
The bound presented in Theorem 4.1 takes the form $\tilde{\mathcal{O}}(D^{2}pd^{2}/(N(T-p)))$. Importantly, the dependencies on the underlying Markov chain are only through $D$ and $\ln\kappa$, which do not have a direct dependency on the mixing time. The dependency on $D$ is merely an operator norm upper bound and does not diverge as the mixing time grows to $\infty$. Similarly, $\ln\kappa$ is logarithmic in $T$ for systems of interest, as we discuss below.
@@
We now explain the behavior of $\kappa$ defined in Definition 3.3. First, by Lemma C.3, we have that $\sigma_{\min}(L_{\star})\geqslant\frac{1}{D+1}$ and, thus, it is sufficient to upper bound
@@
to control $\kappa$. Equation 10 implies that if the noise at step $i$ contributes to step $j$, as measured by $L_{\star}^{(i,j)}$, at a polynomial rate in $\left(j-i\right)$, then $\kappa$ grows at most polynomially in $T$. For such a $\kappa$, the resulting dependency on $T$ is of order $\ln T$ and mild. Instead, if it is exponential in $\left(j-i\right)$, then $\ln\kappa$ grows linearly in $T$ and the dependency on $T$ cancels out in the rate.
We use the quantity $\zeta(T)$ to define strictly stable, marginally stable and explosive systems:
@@
strictly stable if $\quad\zeta(T)=\mathcal{O}(\rho^{T})\quad\text{for some}\quad\rho<1\,,$
marginally stable if $\quad\zeta(T)=\mathcal{O}(T^{k})\quad\text{for some}\quad k\in\mathbb{N}\,,$
explosive if $\quad\zeta(T)=\mathcal{O}(\rho^{T})\quad\text{for some}\quad\rho>1\,.$
Definition 5.1 is similar to the notions of strictly stable, marginally stable and explosive systems considered in (Simchowitz et al., 2018; Sarkar & Rakhlin, 2019) for $p=1$. Let $\rho(A^{\star})\coloneqq\lambda_{\max}(A^{\star})$ be the spectral radius of $A^{\star}$ and $V\Lambda V^{-1}$ be the Jordan normal form of $A^{\star}$. Then,
@@
In particular, for strictly stable systems of Simchowitz et al. (2018); Sarkar & Rakhlin (2019) with $\rho<1$, $\zeta(T)=\mathcal{O}(\rho^{T})$. For marginally stable systems of Sarkar & Rakhlin (2019) with $\rho<1+\frac{\gamma}{T}$ with some constant $\gamma>0$, $\rho(A^{\star})^{j-i}\leqslant e^{\gamma}$ and $\zeta(T)=\mathcal{O}(T^{k})$ for some fixed $k$ that depends on the largest Jordan block of $A^{\star}$. For explosive systems of Sarkar & Rakhlin (2019) with $\rho>1$, $\zeta(T)=\mathcal{O}(\rho^{T})$. Thus, Definition 5.1 provides a general categorization of the systems based on the growth of $\zeta(T)$ in $p>1$ case. Furthermore, our analysis yields sharp rates for strictly stable and marginally stable systems previously considered only in the $p=1$ setting.
@@
Our analysis is based on the assumption that the diameter $D$ of the search space is bounded and, hence, not directly applicable to the OLS estimator in Equation 4. However, Corollary C.1 in Section C.1 extends the results of Theorems 4.1, 4.2 and 4.3 to minimizers without a constraint on the diameter of the search space. This extension does not change the rates but requires the additional assumption that $NT=\tilde{\Omega}(p^{\prime 2}dr)$. [^1] We use the convention that $r=d,p^{\prime}=p$ for Theorem 4.1. In the case of Theorem 4.1, this corresponds to a result for the OLS estimator, but necessitating a number of samples quadratic in context length. Below, we comment on why the diameter restrictions is required when $NT\ll p^{\prime 2}dr$.
As mentioned earlier in comparison with (Simchowitz et al., 2018; Sarkar & Rakhlin, 2019), the simple OLS factorization in the $p=1$ case does not generalize to the $p>1$ and the data matrix has a Toeplitz structure that is more difficult to control. In order to deal with these issues, as explained in the sketch of proof in Appendix A, we rely on techniques from empirical process theory. These techniques are applied to quantify the probability of the event in Equation 13, which holds for any empirical risk minimizer of the square loss. This leads us to the study of the concentration of the martingales defined in Equation 15 around their predictable variation, which is a key step in our analysis. A uniform concentration is possible only if there is a uniform lower bound on the variations of the martingales, which can be achieved using a set of well-behaved matrices $\|M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}}\|_{F}/\|M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}}\|_{\rm op}$. In order to translate these conditions on the design matrices without additional dimensional dependencies, we introduce the operator norm constraint.
@@
The misspecification result in Theorem 4.3 requires the additional assumption given in Assumption 3.5. In Remark C.1, we show that a good upper bound on $D^{\prime}$ is possible when $D<1$, i.e., the system is strictly stable, by using the bound $\|L_{\star}\|_{\rm op}\leqslant 1/(1-D)$. However, misspecification results are not, a priori, applicable to marginally stable systems, which limits the practical applicability of our results. We leave the investigation of misspecification results for marginally stable systems for future work.
Practical implications.
The rates obtained in Section 4 hold true if the empirical risk minimizer ${\boldsymbol{\hat{A}}}$ is replaced by any estimate ${\boldsymbol{\tilde{A}}}$ that verifies
$\mathcal{L}({\boldsymbol{\tilde{A}}})\leqslant\mathcal{L}({\boldsymbol{A}}^{\star})\,.$ (11)
The training error for the ground truth ${\boldsymbol{A}}^{\star}$ concentrates around $\sigma^{2}d$ with a rate of $\mathcal{O}(1/\sqrt{NT})$, which is identical to that in the i.i.d. setting. Hence, any algorithm that optimizes the training error below the threshold $\sigma^{2}d$ achieves the rates in Section 4.
Moreover, in Appendix D, we show that the rates extend to approximate minimizers, i.e., any estimate ${\boldsymbol{\tilde{A}}}$ that verifies
$\mathcal{L}({\boldsymbol{\tilde{A}}})\leqslant\mathcal{L}({\boldsymbol{\hat{A}}})+\epsilon_{\mathrm{tr}}\,,\quad\text{where}\quad\epsilon_{\mathrm{tr}}=\tilde{\mathcal{O}}\left(\frac{p^{\prime}dr}{NT}\right)\,,$ (12)
where $\epsilon_{\mathrm{tr}}$ is the surplus training error of the estimate ${\boldsymbol{\tilde{A}}}$ over the empirical risk minimizer ${\boldsymbol{\hat{A}}}$.
Estimates satisfying equations (11) or (12) are computationally tractable in practice. This implies that practitioners can determine the required number of samples, or $N$ and $T$, for estimating the system parameters up to a fixed precision by using the rates in Section 4.
6 Conclusion
In this work, we extend non-asymptotic linear system identification theory and derive upper bounds on the sample complexity of learning long-context linear autoregressive models. Our bounds improve upon the existing arguments specific to first-order systems by employing a uniform concentration argument over prediction differences. We further establish improved statistical rates when learning under a low-rank assumption. Finally, we show that even with long or unbounded generative contexts, misspecification still allows the estimation of the matrices with a reduced sample complexity for stable systems.
While this work makes significant progress in non-asymptotic linear system identification theory, several technical questions remain open for further investigation. Can the OLS operator norm be coarsely controlled to derive rates for unconstrained OLS in the $NT=\Omega(pdr)$ regime? Is it possible to find efficient algorithms that would benefit from low-rank assumptions? Lastly, can misspecification be beneficial for marginally stable systems?
subsubsectionAcknowledgments This project was supported by the Swiss National Science Foundation (grant number 212111). This work was partially funded by an unrestricted gift from Google.
Appendix A Sketch of Proof
We provide a sketch of proof for Theorem 4.1. The proofs of Theorems 4.2 and 4.3 are similar and can be found in Appendix C. In the following, $\Delta_{{\boldsymbol{A}}}$ is a shorthand for $M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}}$ and $E\in\mathbb{R}^{Td\times N}$ is the matrix that collects the noise concatenated over time, as explained in Definition C.1.
@@ Appendix A Sketch of Proof
$\mathcal{E}\coloneqq{{\left\{\forall{\boldsymbol{A}}\in\mathcal{A}^{\prime}(D):{{\left\|\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right\|}}^{2}\geqslant 2\operatorname{Tr}\left(E^{\top}L_{\star}^{\top}\Delta_{{\boldsymbol{\hat{A}}}}^{\top}E\right)\right\}}}\,,$ (14)
@@
Fix a ${\boldsymbol{A}}\in\mathcal{A}^{\prime}(D)$ and study the martingale series defined by the difference sequences
$d_{t,i}^{(n)}=\left(\left({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}\right)X_{t}^{(n)}\right)_{i}\left(\xi_{t}^{(n)}\right)_{i}\,,$ (15)
@@
$Y_{{\boldsymbol{A}}}=\sum_{n,t,i}d_{t,i}^{(n)}=\sum_{n,t}\left\langle\left({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}\right)X_{t}^{(n)},\xi_{t}^{(n)}\right\rangle=\operatorname{Tr}\left(E^{\top}L_{\star}^{\top}\Delta_{{\boldsymbol{A}}}^{\top}E\right)\,,$
and the predictable quadratic variation of the series is
$W_{{\boldsymbol{A}}}=\sum_{n,t,i}\mathbb{E}_{\left(\xi_{t}^{(n)}\right)_{i}}\left[\left(d_{t,i}^{(n)}\right)^{2}\right]=\sum_{n,t}{{\left\|\left({\boldsymbol{A}}^{\star}-{\boldsymbol{A}}\right)X_{t}^{(n)}\right\|}}^{2}=\sigma^{2}{{\left\|\Delta_{{\boldsymbol{A}}}L_{\star}E\right\|}}^{2}\,.$
The condition in Equation 14 requires that the predictable quadratic variation $W_{{\boldsymbol{A}}}$ is uniformly bounded by a multiple of the sum of the differences $Y_{{\boldsymbol{A}}}$, i.e., =$\mathcal{E}=\left\{\forall{\boldsymbol{A}}\in\mathcal{A}^{\prime}(D):W_{{\boldsymbol{A}}}\leqslant 2\sigma^{2}Y_{{\boldsymbol{A}}}\right\}.$
@@
$\begin{split}\mathbb{P}\left(Y_{\boldsymbol{A}}\geqslant r_{Y}\,,W_{\boldsymbol{A}}^{R}\leqslant r_{W}\right)&\leqslant\exp\left(-\frac{r_{Y}^{2}/2}{r_{W}+Rr_{Y}}\right)\,,\end{split}$ (16)
where $W_{{\boldsymbol{A}}}^{R}=W_{\boldsymbol{A}}+\sum_{n,t,i}\mathds{1}_{d_{t,i}^{(n)}>R}\left(d_{t,i}^{(n)}\right)^{2}$ and $r_{Y},r_{W},R>0$ are arbitrary constants. As the noise is sub-Gaussian, it is possible to upper bound $W_{{\boldsymbol{A}}}^{R}$ with $W_{{\boldsymbol{A}}}$: